Anna has 10 more sheets of paper than Mary. Mary has 7 more sheets of paper than Caroline.They decide to split all of the paper between them equally. How many less sheets of paper would Anna have?
Answer by macston(5194)

.
C=Caroline's sheets; M=Mary's sheets=C+7; A=Anna's sheets=M+10=C+17
.
Total sheets=C+M+A
Total sheets=C+(C+7)+(C+17)=3C+24
.
Equal split=(3C+24)/3=C+8
Each would have 8 more sheets than Caroline originally had.
Since Anna originally had 17 more than Caroline originally had,
she would have 9 fewer sheets.
.
ANSWER: Anna would have 9 fewer sheets.
.